OpenStudy (anonymous):

express the given function h as a composition of f and g so that h(x)=(f of g)(x), where one of the functions is 6x-8 h(x) =(6x-8)^9 can anyone help with this? I'm thrown off.

OpenStudy (anonymous):

\[ f(x) = x^9\\ g(x) = 6 x -8\\ f(g(x)) = f( 6x -8)= (6x -8)^9 \]